OpenStudy (anonymous):

The lengths of fish of a certain type have a normal distribution with mean 38 cm. It is found that 5% of the fish are longer than 50 cm. (i) Find the standard deviation.

OpenStudy (anonymous):

\[z= {d - \mu \over \sigma} \]

OpenStudy (callisto):

Oh... mean + 2sd = 95% of the data 38 + 2sd = 50 2sd = 12 sd = 6 Not sure

OpenStudy (anonymous):

\[p(|z|>{50-38 \over \sigma})=0.05???\]

OpenStudy (anonymous):

\[{50-38 \over \sigma}=0.645\]??

OpenStudy (anonymous):

\[{12 \over \sigma }= 0.645\]??

OpenStudy (anonymous):

\[\sigma = 18.6?\]

OpenStudy (callisto):

For a normal distribution curve, the 95% of the data is = mean + 2 s.d. That means 50 = 38 +2s.d.

OpenStudy (anonymous):

I thought that was the formula I had to use though? I mean the one I wrote?

OpenStudy (anonymous):

|z|>0.05= |z|<0.95= 0.645

OpenStudy (callisto):

I'm not sure. But I think mine is better :P (Just kidding)

OpenStudy (anonymous):

\[this --\sigma-- stands~ for~ standard ~deviation~ though~righ.t.?\]

OpenStudy (callisto):

yes, σ = SD

OpenStudy (anonymous):

Ok. Thanks for helping... but I think I'll stick with my interpretation :p :D

OpenStudy (anonymous):

Thanks for the pic aswell!

OpenStudy (callisto):

Hmm, the pic is for reference, trying to use it to explain my idea. You're welcome

OpenStudy (anonymous):

I found the answer, and I made a little mistake. It's 1.645 instead of 0.645, so the answer is 7.29